A ship is heading due norhtwest at 15 mph in a river that flows east at 7 mph. What is the magnitude and direction of the ship's velocity relative to the Earth's surface?
Answer by stanbon(75887)

A ship is heading due northwest at 15 mph in a river that flows east at 7 mph. What is the magnitude and direction of the ship's velocity relative to the Earth's surface?
-----------------------
Draw the picture on a coordinate system.
The ship vector is heading 135 degrees with magnitude 15
The water vector is heading east with magnitude 7.
----------
The resultant vector is:
x-coordinate: 15cos(135)+7cos(0) = -3.6066...
y-coordinate: 15sin(135)+7sin(0) = 10.6066
--------------------
magnitude = sqrt(-3.6066^2 + 10.6066^2)= 9.9746 mph
direction = arctan(10.6066/-3.6066)=108.78 degrees
=======================================
Cheers,
Stan H.